OpenStudy (anonymous):

At the grocery mart, a 575g can of green beans cost $0.39. How much will an 810 g can of green beans cost?

OpenStudy (texaschic101):

set up a proportion... 575/.39 = 810/x (575g to $.39 = 810 to $x) cross multiply (575)(x) = (.39)(810) 575x = 315.9 x = 315.9/575 x = 0.549 rounds to 55 cents